# The New Science of Unidentified Aerospace-Undersea Phenomena (UAP)

### v1

@@ 1 Introduction
Although people are generally aware that there have been several programs run by the US government to study UFOs (Unidentified Flying Objects) and UAP, it is not generally known precisely what was done and, more importantly, precisely how little was done. Moreover, it is generally believed that serious scientists, academics, and especially astronomers do not believe in the reality of UFOs, have not seen them, and certainly do not study them. By going into some detail on efforts to study UAP in the late 1940s and early 1950s, we learn that some prominent and serious scientists have been involved, that they have witnessed UFOs of many types, and that they took the matter quite seriously. Such details add context Sturrock’s (Sec. A.14) 1977 survey of members of the American Astronomical Society [43] in which he found that out of the 1356 respondents, 23% stated that UFOs should certainly be studied and 30% said that UFOs should probably be studied as opposed to the 17% saying that UFOs should probably not be studied and 3% that said that UFOs should certainly not be studied. Moreover, Sturrock found that 62 of the respondents had either witnessed or obtained an instrumental record of something that they could not explain and that of those responding witnesses, 63% of them were night-sky observers.
@@
Since many countries, research groups, and individuals have participated in the scientific study of UAP for almost a century, there are several ways in which the information in this article could be presented. Unfortunately, there is no single presentation style that appears to be optimal. In some situations the chronology of events and efforts is central, suggesting that a presentation based on chronology is obligatory. However, at other times, once efforts became more focused, it appears to be best to separate out efforts based on the nations, research groups, or the individuals involved. As a result, the presentation in this paper is mixed to a great degree. At some times, especially early on and during wartime, the interactions among different nations dominated the efforts to understand these phenomena. For this reason, the paper begins with a chronological, or historical, treatment which, as efforts shift to become more isolated, gives way to a presentation based on the nations, groups, or individuals involved. An appreciation of both of these differing perspectives is necessary to understand the importance and impact of efforts to study UAP in the past century.
@@ 3 Government Efforts to Study UAP
This paper focuses on the historical efforts to scientifically study UAP. Recent articles by Ailleris provide excellent summaries of such efforts [63, 64]. For those interested in the history of the role of government in this rich and involved topic, we recommend the detailed text by Swords et al. [65], as the governmental aspects of our exposition are guided by their work. As much of the relevant information exists in the gray (non-academic, non-peer-reviewed) literature, it is important to summarize relevant details, if even to dispel some of the common misconceptions about UAP.
@@ 3.3.2 The Wave of 1947
The first prominent encounter was Pilot Kenneth Arnold’s sighting of a long string of nine ‘‘circular type’’ objects near Mount Rainier in Washington on June 24, 1947 [82] (Figure 2A), which was independently corroborated by several other witnesses [79]. This event brought the term ‘‘flying saucer’’ into our lexicon [83]. Arnold estimated the size of the objects to be about 100 feet in diameter, and he used the big sweep 24-hour clock on his instrument panel to time their 1 minute 42 second passage from Mount Rainier to Mount Adams, a distance of about 50 miles. Based on this, he estimated a speed of about 1,760 miles an hour with an uncertainty of a couple hundred miles an hour. This is significant because this estimated speed is almost three times the world airspeed record of 623.62 mph at the time [84, p. 274].
@@ 3.3.5 Green Fireballs and Project Twinkle
It is doubtful that the green fireballs were meteors for another important reason. Meteor storms are global events, which occur when the Earth, in its orbit, passes through an orbiting cometary or asteroid debris cloud. As such, if the green fireballs were meteors, then they should have been observed globally — certainly not limited to a handful of sensitive nuclear sites in New Mexico, and the meteors would have only lasted a few days — not months.
While Project Twinkle was conceived as a data collection effort, it was both underfunded and haphazard in its design. Moving the single camera used in the study meant that the camera kept missing the phenomenon. As a result, it provided virtually no useful information about the phenomena it was intended to study.
The UFO activity at both White Sands Missile Base and Los Alamos in New Mexico resulted in the formation of informal groups of scientists, in both places, to study the phenomena [65]. One member wrote: [103]
@@
To this day, it is not clear what the green fireballs were or if they had any relation to the cylinders, discs, or spheres that had been widely observed. As such, they are an example of the fact that UAP are a class of unknown phenomena, and not any single thing [105]. For this reason, the instruments used to study them need to be sufficiently diverse to be able to provide useful information about a wide class of phenomena.
Moreover, it is a widely held belief that scientists, astronomers in particular, do not witness UFOs or UAP. The events and efforts discussed in this section make it clear that this is not only untrue, but many of the respected scientists who witnessed these phenomena made serious attempts to study them. In addition to illustrating both the seriousness and pervasiveness of these phenomena, knowledge of these events is critical, as they inform us as to what types of equipment have been useful in studying UAP.
Figure 3: An illustration of the disk observed by five adults while boating on the Rogue River in Oregon in May 1949. The observers included two researchers who worked for the National Advisory Committee for Aeronautics (NACA); one of whom worked at the Supersonic Wind Tunnel at Moffett Field, California, and the other who was a draftsman; both of whom were familiar with aerodynamic design. Illustration from Project Blue Book Special Report 14, page 86 [106, 107][108, pp. 863-879, 890-3].
@@ 3.3.6 Project Grudge
To highlight the Air Force’s hostility to UFO cases, it is worth discussing a sighting that occurred in May 1949 on the Rogue River in Oregon [106, 107][108, pp. 863-879, 890-3][65]. A group of six adults were on vacation in a boat on the river when at about 5pm (it was still daylight) they observed a round object that looked like a mirror on edge. The witnesses had binoculars that were passed around so that they all observed the object well. The object had a diameter of about 25–30 feet, was flat on the bottom with an edge like a coin, and had a slightly convex top with what appeared to be a fin-like protrusion on one end (Fig. 3). Two of the witnesses worked for the National Advisory Committee for Aeronautics (NACA, which was the predecessor of NASA). Mr. Don Heaphy was a mechanic at the $1\times 3$ foot Supersonic Wind Tunnel at the Ames Laboratory at Moffett Field, California, and the other was Mr. Gilbert Rivera, a draftsman in the Drafting Section of NACA at Ames Laboratory [106]. As such, both witnesses were uniquely familiar with aerodynamic designs. The object was observed to accelerate from a ‘‘leisurely pace’’ to jet speeds, but there was no jet stream or contrails, and the object made no noise. At one point the object just turned, but the observers noted that it was strange how the object did not need to bank to accomplish the turn. The project report card summarized the event, stating: ‘‘No data presented to indicate object could have NOT been an aircraft. Conclusion: AIRCRAFT’’ [65].
@@ 3.3.7 Project Blue Book
After thousands of reports had been collected, the Condon report, which has been highly criticized, concluded that UFOs did not constitute a direct threat to national security and that the scientific study of UFOs was unlikely to be productive [121]. The Air Force used this conclusion to close Project Blue Book in 1969.
@@ 3.3.9 The Pocantico Meeting
Sturrock described the Review Panel: [127, p. 113]
@@
The Scientific Steering Committee consisted of homas E. Holzer, Robert Jahn, David E. Pritchard, Harold E. Puthoff, Yervant Terzian, and Charles R. Tolbert [127, p.xi]. Sturrock thanked the following scientists (the investigators): [^2] The astute reader will note that Drs. Haines, Rodeghier, Strand, Swords, and Vallée are all coauthors of this present work. Richard F. Haines (Sec. A.4), Illobrand Von Ludwiger, Mark Rodeghier, John F. Schuessler, Erling Strand, Michael D. Swords, Jacques F. Vallée (Sec. A.16), and Jean-Jacques Velasco for providing a crash-course on the UFO topic, and the following scientists: Von R. Eshleman, Thomas E. Holzer, J. R. (Randy) Jokipii, François Louange, H. J. (Jay) Melosh, James J. Papike, Guenther Reitz, Charles R. Tolbert, and Bernard Vayret for serving on the Review Panel and making a ‘‘brave attempt to assess the topic in that very short time’’ [127, p.xi].
Sturrock described the meeting: [127, p.114]
@@
The Review Panel summarized their conclusions and recommendations: [127, pp. 120–122]
@@
The panel concluded with a list of recommendations [127, p. 122], which include:
@@ 3.3.10 United States: The AATIP Program
3.4 Canada: Project Magnet
@@ 3.5 France
Another interesting event that occurred in 1974 was the dissemination of recommendations by members of the Institut des Hautes Études de Défense Nationale (IHEDN) proposing the creation of a UFO study office. These first recommendations, disseminated under the direction of General Blanchard [145], were confirmed by a report in 1977, disseminated by the Association of Auditors, under the leadership of General Richard [166][167, p.300].
@@
Following the departure of Claude Poher in 1978, Alain Esterle took the helm of GEPAN. Under his direction, the organization published numerous technical documents [168, 169, 170, 171, 172, 173, 174, 175, 176], information notes [177, 178, 179] and working documents [180, 181, 182]. Alain Esterle left the group in 1983 and was replaced by Jean-Jacques Velasco, who led it for 20 years. Under his direction, GEPAN was renamed the Atmospheric Reentry Phenomena Expertise Service (SEPRA) in 1988 [183]. The structure operated with a network of investigators making it possible to meet witnesses after they have given testimony to the gendarmerie, and after an initial field investigation by the gendarmes. The GEPAN and then SEPRA investigators had specific equipment, including the SIMOVNI developed in part by Mr. Velasco, making it possible to simulate in the field what the witness observed, as well as an inclinometer to measure the elevation of the object in relation to the visual field of the witness. Although GEPAN enjoyed a certain level of independence [184], CNES decided to withdraw the organization’s ability to conduct further scientific studies, after some failed to achieve their goals [185, p.98–99]. It seemed that all GEPAN could do now was collect data. The scientific programs were now ordered by the CNES Scientific Programs Committee. However, even though GEPAN / SEPRA could not finance them, it could help other research actions [186].
@@ 3.6 Russia
The Soviet Navy was also involved from 1977 (07/10/1977) in research on abnormal phenomena. It issued guidelines for investigating and gathering information on UFOs to fleets and flotillas, by order of Admiral Smironov, Deputy Chief of Staff of the Navy. This was then extended on January 20, 1978 by Vice-Admiral Ivanov, Chief of the Intelligence Directorate of the Navy, who gave instructions to the Oceanographic Commission, with the assistance of V. G. Ajaja. According to the KGB, mention is made of cases of material collection with particular properties on certain sites such as Dalnegorsk. This work seems to have stopped or to have been greatly slowed down after 1990.

### v2

@@ 1 Introduction
Although people are generally aware that there have been several programs run by the US government to study UFOs (Unidentified Flying Objects) and UAP, it is not generally known precisely what was done and, more importantly, precisely how little was done. Moreover, it is generally believed that serious scientists, academics, and especially astronomers do not believe in the reality of UFOs, have not seen them, and certainly do not study them. By going into some detail on efforts to study UAP in the late 1940s and early 1950s, we learn that some prominent and serious scientists have been involved, that they have witnessed UFOs of many types, and that they took the matter quite seriously. Such details add context Sturrock’s (Sec. A.15) 1977 survey of members of the American Astronomical Society [43] in which he found that out of the 1356 respondents, 23% stated that UFOs should certainly be studied and 30% said that UFOs should probably be studied as opposed to 17% saying that UFOs should probably not be studied and 3% that said that UFOs should certainly not be studied. Moreover, Sturrock found that 62 of the respondents had witnessed or obtained an instrumental record of something that they could not explain, and that of the respondent witnesses, 63% of them were night-sky observers.
@@
Since many countries, research groups, and individuals have participated in the scientific study of UAP for almost a century, there are several ways in which the information in this article could be presented. Unfortunately, there is no single presentation style that appears to be optimal. In some situations, the chronology of events and efforts is central, suggesting that a presentation based on chronology is obligatory. However, at other times, once efforts became more focused, it appears to be best to separate out efforts based on the nations, research groups, or the individuals involved. As a result, the presentation in this paper is mixed to a great degree. At some times, especially early on and during wartime, the interactions among different nations dominated the efforts to understand these phenomena. For this reason, the paper begins with a chronological, or historical, treatment which, as efforts shift to become more isolated, gives way to a presentation based on the nations, groups, or individuals involved. An appreciation of both of these differing perspectives is necessary to understand the importance and impact of efforts to study UAP in the past century.
@@ 3 Government Efforts to Study UAP
This paper focuses on the historical efforts to scientifically study UAP. Recent articles by Ailleris [63, 64], Stahlman [65], Villaroel & Krisciunas [66], and Watters et al. [62] provide excellent summaries of such efforts. For those interested in the history of the role of government in this rich and involved topic, we recommend the detailed text by Swords et al. [67], as the governmental aspects of our exposition are guided by their work. As much of the relevant information exists in the gray (non-academic, non-peer-reviewed) literature, it is important to summarize relevant details, if even to dispel some of the common misconceptions about UAP.
@@ 3.3.2 The Wave of 1947
The first prominent encounter was Pilot Kenneth Arnold’s sighting of a long string of nine ‘‘circular type’’ objects near Mount Rainier in Washington on June 24, 1947 [84] (Figure 2A), which was independently corroborated by several other witnesses [81]. This event brought the term ‘‘flying saucer’’ into our lexicon [85]. Arnold estimated the size of the objects to be about 100 feet in diameter, and he used the big sweep 24-hour clock on his instrument panel to time their 1 minute 42 second passage from Mount Rainier to Mount Adams, a distance of about 50 miles. Based on this, he estimated a speed of about 1,760 miles an hour with an uncertainty of a couple hundred miles an hour. This is significant because this estimated speed is almost three times the world airspeed record of 623.62 mph at the time [86, p. 274].
@@ 3.3.5 Green Fireballs and Project Twinkle
It is doubtful that the green fireballs were meteors for another important reason. Meteor storms are global events that occur when the Earth, in its orbit, passes through an orbiting cometary or asteroid debris cloud. As such, if the green fireballs were meteors, then they should have been observed worldwide — certainly not limited to a handful of sensitive nuclear sites in New Mexico, and the meteors would have only lasted a few days — not months.
Although Project Twinkle was conceived as a data collection effort, it was both underfunded and haphazard in its design. Moving the single camera used in the study meant that the camera kept missing the phenomenon. As a result, it provided virtually no useful information about the phenomena it was intended to study.
The UFO activity at both the White Sands Missile Base and Los Alamos in New Mexico resulted in the formation of informal groups of scientists, in both places, to study the phenomena [67]. One member wrote: [105]
@@
To this day, it is not clear what the green fireballs were or if they had any relation to the cylinders, discs, or spheres that had been widely observed. As such, they are an example of the fact that UAPs are a class of unknown phenomena, and not a single thing [107]. For this reason, the instruments used to study them need to be sufficiently diverse to be able to provide useful information about a wide class of phenomena.
Moreover, it is a widely held belief that scientists, astronomers in particular, do not witness UFOs or UAP. The events and efforts discussed in this section make it clear that this is not only untrue, but many of the respected scientists who witnessed these phenomena made serious attempts to study them. In addition to illustrating both the seriousness and pervasiveness of these phenomena, knowledge of these events is critical, as they inform us of what types of equipment have been useful in studying UAP.
Figure 3: An illustration of the disk observed by five adults while boating on the Rogue River in Oregon in May 1949. The observers included two researchers who worked for the National Advisory Committee for Aeronautics (NACA); one of whom worked at the Supersonic Wind Tunnel at Moffett Field, California, and the other who was a draftsman; both of whom were familiar with aerodynamic design. Illustration from Project Blue Book Special Report 14, page 86 [108, 109][110, pp. 863-879, 890-3].
@@ 3.3.6 Project Grudge
To highlight the Air Force’s hostility to UFO cases, it is worth discussing a sighting that occurred in May 1949 on the Rogue River in Oregon [108, 109][110, pp. 863-879, 890-3][67]. A group of six adults were on vacation in a boat on the river when at about 5pm (it was still daylight) they observed a round object that looked like a mirror on edge. The witnesses had binoculars that were passed around so that they all observed the object well. The object had a diameter of about 25–30 feet, was flat on the bottom with an edge like a coin, and had a slightly convex top with what appeared to be a fin-like protrusion on one end (Fig. 3). Two of the witnesses worked for the National Advisory Committee for Aeronautics (NACA, which was the predecessor of NASA). Mr. Don Heaphy was a mechanic at the $1\times 3$ foot Supersonic Wind Tunnel at the Ames Laboratory at Moffett Field, California, and the other was Mr. Gilbert Rivera, a draftsman in the Drafting Section of NACA at Ames Laboratory [108]. As such, both witnesses were uniquely familiar with aerodynamic designs. The object was observed to accelerate from a ‘‘leisurely pace’’ to jet speeds, but there was no jet stream or contrails, and the object made no noise. At one point the object just turned, but the observers noted that it was strange how the object did not need to bank to accomplish the turn. The project report card summarized the event, stating: ‘‘No data presented to indicate object could have NOT been an aircraft. Conclusion: AIRCRAFT’’ [67].
@@ 3.3.7 Project Blue Book
After thousands of reports had been collected, the Condon report [123], which has been highly criticized, concluded that UFOs did not constitute a direct threat to national security and that the scientific study of UFOs was unlikely to be productive [124]. The Air Force used this conclusion to close Project Blue Book in 1969.
@@ 3.3.9 The Pocantico Meeting
Sturrock described the Review Panel: [130, p. 113]
@@
The Scientific Steering Committee consisted of homas E. Holzer, Robert Jahn, David E. Pritchard, Harold E. Puthoff, Yervant Terzian, and Charles R. Tolbert [130, p.xi]. Sturrock thanked the following scientists (the investigators): [^2] The astute reader will note that Drs. Haines, Rodeghier, Strand, Swords, and Vallée are all coauthors of this present work. Richard F. Haines (Sec. A.5), Illobrand Von Ludwiger, Mark Rodeghier, John F. Schuessler, Erling Strand, Michael D. Swords, Jacques F. Vallée (Sec. A.17), and Jean-Jacques Velasco for providing a crash-course on the UFO topic, and the following scientists: Von R. Eshleman, Thomas E. Holzer, J. R. (Randy) Jokipii, François Louange, H. J. (Jay) Melosh, James J. Papike, Guenther Reitz, Charles R. Tolbert, and Bernard Vayret for serving on the Review Panel and making a ‘‘brave attempt to assess the topic in that very short time’’ [130, p.xi].
Sturrock described the meeting: [130, p.114]
@@
The Review Panel summarized their conclusions and recommendations: [130, pp. 120–122]
@@
The panel concluded with a list of recommendations [130, p. 122], which include:
@@ 3.3.10 United States: The AATIP Program
3.4 Canada
3.4.1 Project Magnet
@@
3.4.2 Sky Canada Project
The Sky Canada Project, led by the Office of the Chief Science Advisor (OCSA) of Canada, was launched in the Fall of 2022 to study and evaluate how UAP sightings are managed in Canada [144]. A preliminary report was published in January 2025 [145] with a full report expected to be publsihed in March 2025.
Sky Canada works to collect information from relevant federal departments and agencies, experts, non-government organizations, and a number of countries, including the G7 nations and members of the Five Eyes. These federal departments include the Canadian Coast Guard, the Canadian Nuclear Safety Commission, the Canadian Space Agency, the Department of National Defence (DND), the Defence Research and Development Canada (DRDC), GEIPAN (France), the Library and Archives Canada (LAC), the Meteorological Service of Canada (MSC) at Environment and Climate Change Canada (ECCC), MUFON Canada, the National Research Council of Canada (NRC), NAV CANADA, the Ontario Provincial Police (OPP), the Royal Canadian Mounted Police (RCMP), Sûreté du Québec (SQ), and Transport Canada [145].
The preliminary report notes that Canadians report 600 to 1,000 sightings annually, with one in four respondents claiming to have personally witnessed a UAP in their lifetime. Sky Canada recommends that 1) data on UAPs be made publicly available to support research, 2) periodic surveys be conducted to gauge public perception and to improve services, 3) participatory science programs should be developed to enable volunteer involvement in UAP studies, and 4) publicly accessible digital tools should be made available to aid in data collection. Furthermore, Sky Canada recommends developing ‘‘partnerships with international entities dedicated to UAPs, such as AARO and NASA (U.S.A.), GEIPAN (France) and SEFAA (Chile), to share data, methodologies and best practices in UAP research and investigation’’ [144].
The Sky Canada preliminary report concludes by stating: [145]
Adopting a science-based, collaborative approach will help address public concerns, demystify UAPs, and potentially reveal valuable insights into aerial phenomena that are currently unexplained.
@@ 3.5 France
Another interesting event that occurred in 1974 was the dissemination of recommendations by members of the Institut des Hautes Études de Défense Nationale (IHEDN) proposing the creation of a UFO study office. These first recommendations, disseminated under the direction of General Blanchard [148], were confirmed by a report in 1977, disseminated by the Association of Auditors, under the leadership of General Richard [169][170, p.300].
@@
Following the departure of Claude Poher in 1978, Alain Esterle took the helm of GEPAN. Under his direction, the organization published numerous technical documents [171, 172, 173, 174, 175, 176, 177, 178, 179], information notes [180, 181, 182] and working documents [183, 184, 185]. Alain Esterle left the group in 1983 and was replaced by Jean-Jacques Velasco, who led it for 20 years. Under his direction, GEPAN was renamed the Atmospheric Reentry Phenomena Expertise Service (SEPRA) in 1988 [186]. The structure operated with a network of investigators making it possible to meet witnesses after they have given testimony to the gendarmerie, and after an initial field investigation by the gendarmes. The GEPAN and then SEPRA investigators had specific equipment, including the SIMOVNI developed in part by Mr. Velasco, making it possible to simulate in the field what the witness observed, as well as an inclinometer to measure the elevation of the object in relation to the visual field of the witness. Although GEPAN enjoyed a certain level of independence [187], CNES decided to withdraw the organization’s ability to conduct further scientific studies, after some failed to achieve their goals [188, p.98–99]. It seemed that all GEPAN could do now was collect data. The scientific programs were now ordered by the CNES Scientific Programs Committee. However, even though GEPAN / SEPRA could not finance them, it could help other research actions [189].
@@ 3.6 Russia
The Soviet Navy was also involved from 1977 (07/10/1977) in research on abnormal phenomena. It issued guidelines for investigating and gathering information on UFOs to fleets and flotillas, by order of Admiral Smironov, Deputy Chief of Staff of the Navy. This was then extended on January 20, 1978, by Vice Admiral Ivanov, Chief of the Intelligence Directorate of the Navy, who gave instructions to the Oceanographic Commission, with the assistance of V. G. Ajaja. According to the KGB, mention is made of cases of material collection with particular properties on certain sites such as Dalnegorsk. This work seems to have stopped or to have been greatly slowed down after 1990.